MOST REACTIVE ELEMENT IN THE PERIODIC TABLE ….............................................................................
 
 

Grade:9

5 Answers

Arun
6 years ago
Let the output of both engines be WLet the 1st engine take Q1 heat as input at temperature T1 = 800K and gives out heat Q2 at temperature T the second engine receive Q2 as input and give out Q3 at temperature T3 = 300k to the sink.Work done by first engine W = Q1 – Q2Work done by 2nd engine W = Q2 – Q3Thus, Q1 – Q2 = Q2 – Q3Dividing both sides by Q11 – Q2/Q1 = Q2/Q1 – Q3/Q1=> 1-T/T1 = Q2/ Q1(1- Q3/ Q2)=> 1-T/T1 = Q2/ Q1(1- T3/ T)=> 1-T/T1 = T/T1 (1- T3/ T)=> T1/T – 1 = 1 - T3/ T=> T1/T + T3/ T = 2=> 1/T (T1 + T3) = 2=>T = (T1 + T3)/2=> T = (800+300)/2 = 550 K
Arun
6 years ago
The most reactive metal on the periodic table is francium. However, francium is a man-made element and only minute quantities have been produced, so for all practical purposes, the most reactive metal is cesium. Cesium reacts explosively with water. It is predicted francium would react even more vigorously.
sahil
6 years ago
The most reactive metal on the periodic table is francium but due to its quantity cesium is the answer
praveen kumar
6 years ago
cesium is the most reactive element in the periodic table because of it’s size the nuclear attraction is very very less so that it can lose electrons very easily and can participate in the chemical reaction.
Sayantan Garai
5 years ago
F is the most reactive element.
Cs is the most reactive metal.
F is a non-metal.
You can find the properties of them on IInternet.